x^3*dy/dx – x^2*y = – y^4*cosx


1
Expert's answer
2021-02-24T07:10:56-0500

"x^3\frac{dy}{dx}-x^2y=-y^4\cos{x}"

Solution:

Divide the left and right sides of the equation by "x^2y^4" :

"\frac{x}{y^4}\frac{dy}{dx}-\frac{1}{y^3}=-\frac{\cos{x}}{x^2}"

"\frac{1}{y^3}=z" , "z'=-3\frac{y'}{y^4}" , "\frac{y'}{y^4}=-\frac{z'}{3}" .

"x(-\frac{z'}{3})-z=-\frac{\cos{x}}{x^2}"

"x\frac{z'}{3}+z=\frac{\cos{x}}{x^2}"

"z=uv" , "z'=u'v+v'u"

"x\frac{u'v+v'u}{3}+uv=\frac{\cos{x}}{x^2}"

"x\frac{u'v}{3}+u(\frac x3v'+v)=\frac{\cos{x}}{x^2}"

Let's compose and solve the system:

"\begin{cases}
 \frac x3v'+v=0 \\
 x\frac{u'v}{3}=\frac{\cos{x}}{x^2} 
\end{cases}"

From the first equation:

"\frac{dv}{v}=-3\frac{dx}{x}"

"\ln|v|=-3\ln|x|"

"v=\frac{1}{x^3}"

Substitute "v" into the second equation:

"x\frac{u'}{3x^3}=\frac{\cos{x}}{x^2}"

"u'=3\cos{x}"

"u=3\sin{x}+C"

"z=uv=\frac{1}{x^3}(3\sin{x}+C)"

"\frac{1}{y^3}=\frac{1}{x^3}(3\sin{x}+C)"

"y^3=\frac{x^3}{3\sin{x}+C}"

Answer: "y^3=\frac{x^3}{3\sin{x}+C}" .